Question:
Find the vertex, focus, and directrix of the grph of the equation
16(y+7)=(x+1)^2
(A) vertex (1,-7) focus (-1,11) directrix is y=-5
(B) vertex (-1,-7) focus (-1,-3) directrix is y=-11
(C) vertex (1,-7) focus (1,-3) directrix is y=-3
(D) vertex (-1,-7) focus (1,11) directris is y=3
Thank you in advance for your help!

Answer by mukhopadhyay(490)

This is a function of a parabola., opening upward.
16(y+7)=(x+1)^2
=> (x+1)^2 = 4*4*(y+7)
The vertex is at (-1,-7);
...
The focus is (-1,-3) (same x-coordinate, y-coordinate is 4 units up from the y-coordinate of the vertex)
.....
Directrix is parallel to x-axis and is located 4 units below the vertex
=> Directrix is y = -11
...
So, the correct option for the answer is B
..
When you send a "Thank you" message, would you please indicate the grade level in which such problems are covered?... Thanks